OpenStudy (gibby):

Describe the graph of the function f(x) = x3 − 15x2 + 68x − 96. Include the y-intercept, x-intercepts, and the shape of the graph.

OpenStudy (gibby):

Someone please help

jimthompson5910 (jim_thompson5910):

If you replace every x with 0, what is the result? In other words, what is the value of f(0)?

OpenStudy (gibby):

0.. im not really sure how to do any of it.

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = x^3 - 15x^2 + 68x - 96\] \[\Large f({\color{red}{x}}) = ({\color{red}{x}})^3 - 15({\color{red}{x}})^2 + 68({\color{red}{x}}) - 96\] \[\Large f({\color{red}{0}}) = ({\color{red}{0}})^3 - 15({\color{red}{0}})^2 + 68({\color{red}{0}}) - 96\] \[\Large f(0) = ???\] Use PEMDAS to simplify. Tell me what you get.

OpenStudy (gibby):

-96?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

when x = 0, it leads to y = -96

jimthompson5910 (jim_thompson5910):

so (x,y) = (0,-96) is the y intercept as an ordered pair. This is where it crosses the y axis.

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (gibby):

Yes a little bit. So is that how i would fully graph it? like what do i do for that?

jimthompson5910 (jim_thompson5910):

Is your teacher letting you use a graphing calculator or technology to graph this? or does your teacher want you to do it by hand?

OpenStudy (gibby):

by hand.

jimthompson5910 (jim_thompson5910):

ok so what you'll need to do is generate a table of points

jimthompson5910 (jim_thompson5910):

(0,-96) is one point. Let's find another when x = 1, then \[\Large f(x) = x^3 - 15x^2 + 68x - 96\] \[\Large f({\color{red}{x}}) = ({\color{red}{x}})^3 - 15({\color{red}{x}})^2 + 68({\color{red}{x}}) - 96\] \[\Large f({\color{red}{1}}) = ({\color{red}{1}})^3 - 15({\color{red}{1}})^2 + 68({\color{red}{1}}) - 96\] \[\Large f(1) = ???\] Use PEMDAS to simplify. Tell me what you get.

OpenStudy (gibby):

-43

jimthompson5910 (jim_thompson5910):

close but no

OpenStudy (gibby):

Can you explain to me how this one is done

jimthompson5910 (jim_thompson5910):

Replace every x with 1 to get \[\Large f(x) = x^3 - 15x^2 + 68x - 96\] \[\Large f({\color{red}{x}}) = ({\color{red}{x}})^3 - 15({\color{red}{x}})^2 + 68({\color{red}{x}}) - 96\] \[\Large f({\color{red}{1}}) = ({\color{red}{1}})^3 - 15({\color{red}{1}})^2 + 68({\color{red}{1}}) - 96\] \[\Large f(1) = 1 - 15(1) + 68(1) - 96\] \[\Large f(1) = 1 - 15 + 68 - 96\] \[\Large f(1) = -14 + 68 - 96\] \[\Large f(1) = 54 - 96\] \[\Large f(1) = -42\]

OpenStudy (gibby):

oh wow i was close lol. that makes sense

OpenStudy (gibby):

So whats the next step.

OpenStudy (jalil.h):

jim can u help me plz im in a hurry

jimthompson5910 (jim_thompson5910):

yes you were only off by 1 @Gibby

jimthompson5910 (jim_thompson5910):

so x = 1 leads to y = -42

jimthompson5910 (jim_thompson5910):

(x,y) = (1,-42) is another ordered pair on this graph

jimthompson5910 (jim_thompson5910):

when x = 2, what is the value of y?

OpenStudy (gibby):

What do you mean. Im sorry if I'm being bit complicated. Math is not the subject for me.

jimthompson5910 (jim_thompson5910):

plug in x = 2 (i.e., replace every x with 2). What do you get when you simplify?

OpenStudy (gibby):

-20

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (gibby):

Was i close

jimthompson5910 (jim_thompson5910):

kinda

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (gibby):

ok

OpenStudy (gibby):

-12

jimthompson5910 (jim_thompson5910):

yes so (2,-12) is another point on this curve

jimthompson5910 (jim_thompson5910):

so you just keep doing this to generate multiple points. You will plot the points and then draw a curve through them all. The more points you determine, the more accurate the graph.

OpenStudy (gibby):

Ok. So how many do you suggest i do.

jimthompson5910 (jim_thompson5910):

not sure. At least 5 or so points. That way you can see the general shape and direction and such

OpenStudy (gibby):

So when i graph it i just connect the points because doesn't it have to be rounded. i just feel like i wouldn't get it right.

jimthompson5910 (jim_thompson5910):

yes you would draw a curve through all the points the best you can. Since it's done by hand, it doesn't have to be perfect

OpenStudy (gibby):

So thats all for this problem?

jimthompson5910 (jim_thompson5910):

That's just for the graph. For the x intercepts, you'll have to use a graphing calculator to find them. If you can't use a graphing calculator, then you'll have to make estimates based on your graph

OpenStudy (gibby):

How do i do that by hand?

jimthompson5910 (jim_thompson5910):

you mean find the x intercepts? or do the graph?

OpenStudy (gibby):

the x intercepts

jimthompson5910 (jim_thompson5910):

you would look to see where the graph crosses or touches the x axis

OpenStudy (gibby):

oh i see ok

jimthompson5910 (jim_thompson5910):

since your graph won't be perfect, neither will the x intercept estimates. But that's to be expected

OpenStudy (gibby):

Ok. so is that all i have to do?

jimthompson5910 (jim_thompson5910):

yes that will wrap up the problem

OpenStudy (gibby):

Ok thank you! I really appreciate it!

jimthompson5910 (jim_thompson5910):

glad to be of help